How to solve for H. The equation reads.
S=4LW + 2WH; S=96, L=5, W=4

Answer by nyc_function(2741)

96 = 4(5)(4) + 2(4)H


96 = 80 + 8H


96 - 80 = 8H


16 = 8H


16/8 = H


2 = H

Done!